OpenStudy (anonymous):

y varies jointly with p and q. When p = 5 and q = 2, y = 80. Find y when p = 4 and q = 5.

OpenStudy (agent47):

y=kpq 80=k*5*2=10k k=8

OpenStudy (agent47):

y=8pq y=8*4*5=8*20=160

OpenStudy (anonymous):

thank u :)